OpenStudy (anonymous):

An official major league baseball has a mass of 0.14 kg. A pitcher throws a 40 m/s fastball which is hit by the batter straight back up the middle at a speed of 46 m/s. A) What is the change in momentum of the ball during the collision with the bat? B) If this collision occurs during a time of 0.012 seconds, what is the average force exerted by the bat on the ball?

OpenStudy (shamim):

Given m=0.14kg u=40m/s v=-46m/s

OpenStudy (shamim):

A) Change in momentum=mv-mu=?

OpenStudy (shamim):

B) Force F=(mv-mu)/t=?

OpenStudy (shamim):

Response plz!!!